A jetplane and a tanker plane are 350 miles apart. They head towards each other so the jet can refuel. The jet flies 250 mi/hr faster than the tanker. Determine the speed of each aircarft if they meet in thrity minutes.
Okay, so i realize this is the type of problem where you have to come up with two equations then use the subsitution method to solve for each variable, but I cannot figure out what the two equations would be.
So What I have so far is-
T-Speed of Tanker
J-Speed of Jet
J= T+250
Then i have tried several equations for the second one, but none of them seem right! I am completley stuck! Please help me out if at all possible! Thanks a bunch!

Answer by stanbon(57328)

A jetplane and a tanker plane are 350 miles apart. They head towards each other so the jet can refuel. The jet flies 250 mi/hr faster than the tanker. Determine the speed of each aircraft if they meet in thirty minutes.
--------------------------------------
Rate Equation: distance/time = rate
350 mi/(1/2)hr = 700 mph
-------------------------------
Tankar speed is x ; Jet speed = 250
Equation:
x + (x + 250) = 700
2x = 450
x = 225 mph (tankar speed)
x+50 = 275 mph (jet speed)
-------------------------------
Cheers,
Stan H.